GDP and the Measurement of ALFRED i Progress - Ask FRED 12,000 The accompanying graph depicts the components of quarterly Real GDP in the United States measured in billions of dollars: O 10,000 8,000 Real Personal Consumption Expenditures Vintage: 2017-12 21 Q1 2010 9,917.689 6,000 GDP = C + I + G + NX 4,000 Real Government Consumption Q1 2010 3,084.280 Expenditures and Gross Investment Vintage: 2018-01-23: Real Gross Private Domestic 1 2010 2,012.899 Investment Vintage: 2018-01-23: Net Exports of Goods and Services Q1 2010 -488.704 Vintage: 2018-01-23 2,000 C: Real personal consumption expenditures » I: Real private domestic 2,000 investment 2016 * G: Real government consumption expenditures and gross investment Source: U.S. Bureau of Economic Analysis ·NX: Real net exportsb. In the first quarter of 2010, what share of GDP was made up by consumption? Please enter your answer as a percentage, and round your answer to one place after the decimal. Consumption share of GDP:
Answer #1

In the first quarter, GDP was C + I + G + NX = 9917.689 + 3084.28 + 2012.899 - 488.704 = 14526.16 billion of dollars

Consumption share in GDP = 9917.689 x 100 / 14526.16 = 68.27%